OpenStudy (anonymous):

which of the following will give you an incorrect slope? A. Y2-Y1/(<--over) X2-X1 B. Y2-Y1/(<--over) X1-X2 C. Y1-Y2/(<--over) X1-X2 D. Rise/(<--over) Run

OpenStudy (kidrah69):

\[\frac{ y2-y1 }{ x2-x1 }=slope\] which does not match

OpenStudy (kidrah69):

\[\frac{ rise }{ run }\] also gives slope

OpenStudy (kidrah69):

Are we only allowed to pick ONE?

OpenStudy (anonymous):

yes

OpenStudy (kidrah69):

alright so its saying which does NOT i already eliminated 2 for you can u eliminate the other one? :)

OpenStudy (kidrah69):

Let me clue you in...now that you know the slope formula which of your options switches the X and Y around?

OpenStudy (jdoe0001):

hint: \(\bf slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}}\)

OpenStudy (anonymous):

y2-y1 over x1-x2?

OpenStudy (jdoe0001):

for the slope, yes

OpenStudy (anonymous):

thank you!!!!!!

OpenStudy (anonymous):

thank you!!!!!!